What is the value of x in the equation 4(2x + 12) = 0?
Found 4 solutions by addingup, Alan3354, stanbon, Shana-D:
Answer by addingup(3677)

4(2x + 12) = 0
8x + 48 = 0
8x = -48
x = -48/8
x = -6
Proof:
4(2(-6) + 12)= 0
4(-12 + 12)= 0
4(0) = 0
0 = 0
The answer is correct.

Answer by Alan3354(69443)

What is the value of x in the equation 4(2x + 12) = 0?
--------------
4(2x + 12) = 0
Step 1, divide by 4
Step 2, divide by 2

Answer by stanbon(75887)

4(2x + 12) = 0
Divide both sides by 4 to get:
2x+12 = 0
----
Divide thru by 2 to get:
x + 6 = 0
------
Subtract 6 from both sides to get:
x = -6
---------------------
Cheers,
Stan H.
-------------

Answer by Shana-D(6)

4(2x + 12) = 0
8x + 48 = 0
8x = -48
x = -6